# Out-of-equilibrium Eigenstate Thermalization Hypothesis

### v4

@@ 1 Introduction
One of the central questions is how to extend the ETH framework to describe non-equilibrium dynamics [2, 35, 36, 37, 38, 39, 40, 41]. In this work, we propose a statistical ansatz for the matrix elements of the projector on the initial out-of-equilibrium state $\Psi=\ket{\psi}\bra{\psi}$ | written in the eigenbasis of the Hamiltonian. Notably, the non-equilibrium dynamics are encoded in the correlations between the initial state and the observable’s off-diagonal matrix elements, which we describe in our framework. After introducing the ansatz and verifying its consistency, we discuss its implications for the relaxation dynamics towards equilibrium and numerically verify it in a non-integrable one-dimensional spin chain. The main novelty of this work stems in the interplay between the standard ETH ansatz for the operator and the one that we propose for the projector via the correlations that characterise their matrix elements and that allows us to describe the non-equilibrium dynamics, beyond the steady value.
@@ 2.2 Ansatz
Diagonal pseudo-random $\tilde{R}_{ii}$ also have zero average, but the particular value of their variance may depend on the symmetry class of $\hat{H}$ (e.g. GOE or GUE), as it is the case for $R_{ii}$ in standard ETH [44, 19]. Given that $\Psi_{ii}$ is positive, $(1+R_{ii})\geq 0$. Since $\Psi_{ij}$ is a product of two quasi-random numbers, this implies various constraints on the joint properties of $\tilde{R}_{ij}$. Crucially, these variables are exponentially weakly correlated with $R_{ij}$ of the original ETH ansatz (1),
$\overline{R_{ij}\tilde{R}_{ji}}=g_{A,\Psi}(e^{+}_{ij,},\omega_{ij})\,e^{-S(e^{+}_{ij})/2}\ ,$ (8)
where $g_{A,\Psi}(e^{+}_{ij,},\omega_{ij})$ is an order one smooth function of its variables, which describes the correlations crucial for non-equilibrium dynamics. The existence of such correlations between the off-diagonal products of $\Psi_{ij}A_{ji}$, can be shown by averaging over the (random) phases of the eigenvectors of $\hat{H}$, see [45].
@@
$\overline{|\Psi_{ij}|^{2}}$ $=\frac{e^{-2N\phi(e^{+})}}{Z^{2}}\,e^{-\frac{\phi^{\prime\prime}(e^{+})}{4N}\omega_{ij}^{2}},$ (10b)
@@
where we made explicit the dependence on the system size $N$. Eq.(10b) applies to the states conforming to (4) and it is obtained by expanding $E_{i,j}=E^{+}\pm\omega_{ij}/2$ around $E^{+}$ using the assumption of large deviation. A similar term, $e^{-\frac{\phi^{\prime\prime}(e^{+})}{8N}\omega_{ij}^{2}}$, should appear also in Eq.(10c), however, in the limit of large $N$ this can be neglected because $f_{A}(e,\omega)$ is expected to decay at large frequencies at least as $\exp(-\beta/4\omega_{ij})$ [46] and only close-by eigenvectors contribute to quantum averages.
Finally, we note that the product $\Psi_{ij}A_{ij}$ has large fluctuations compared to the average (10c), see [45].
@@ 2.3 Consistency checks
At the leading order in $N$, this implies, see [45],
@@ 2.4 Implications for the dynamics
By substituting sums with integrals, neglecting spectral correlations, and using the out-of-equilibrium ETH ansatz in Eq.(10b), at the leading order in $N$, one finds
@@
where we used the definition of the energy variance in Eq.(11), $\Delta_{E_{0}}=1/\sqrt{\Phi^{\prime\prime}(E_{0})-S^{\prime\prime}(E_{0})}$ $=\delta_{e_{0}}\sqrt{N}$. Thus the large deviation ansatz in Eq.(6) is consistent with the Gaussian decay, controlled by the energy variance of the initial state, in agreement with the literature on global quenches, see e.g. Ref. [48]. Dynamical behavior, different from Eq.(16), e.g. including an exponential decay, is known to arise from the initial states which are different from Eq.(4) [47, 51, 52, 48, 49, 53], as discussed below.
@@
Similarly to $f_{A}^{2}(\omega)$, which has to decay exponentially at large $\omega$ in $D\geq 2$ (superexponentially in 1D), high-frequency tail of $g_{\Psi,A}(\omega)$ has to be exponentially suppressed for states $\Psi$ associated with local perturbations, see [45].
@@ 3 Numerical results
Figure 1: Out-of-Equilibrium ETH of the fully polarized initial state ${|\psi\rangle=|\downarrow_{z}\dots\downarrow_{z}\rangle}$. (a) The rescaled diagonal ensemble as a function of the energy density for different system sizes $L=12,14,16,18$. In the inset, the initial energy $E_{0}=\langle\psi|\hat{H}|\psi\rangle$ and the energy fluctuations $(\Delta E_{0})^{2}=\langle\psi|\hat{H}^{2}|\psi\rangle-E_{0}^{2}$ are plotted as a function of the inverse of the system size $1/L$. In both panels, the blue dots correspond to individual overlaps for $L=16$. The smoothing parameter is $\tau=4$. In the inset, the numerical fidelity decay up to $L=22$ is compared with the prediction in Eq.(16) (dashed) without fitting parameter.
First, we establish that the initial states (23) are consistent with the ansatz in Eq.(6). In Fig. 1a we plot the diagonal ensemble for different length sizes $L=12,14,16,18$, showing that it obeys the large deviation prediction $\overline{\Psi_{ii}}=\frac{e^{-L\phi(E_{i}/L)}}{Z}$. This is confirmed by the inset, where we plot the scaling of the initial energy $E_{0}=e_{0}L$ and variance $\Delta_{E_{0}}/E_{0}=\delta_{e_{0}}/e_{0}\sqrt{L}$, c.f. Eq.(11). From a fit of the data, we extract the dimensionless values $e_{0}=0.10$, $\delta_{e_{0}}=1.12$. In panel (b), we study the fluctuations of the out-of-equilibrium ETH functions in the frequency domain [cf. Eq.(10b)]. To address the dependence on $L$, we re-scale $\overline{|\Psi_{ij}|^{2}}$ by the diagonal matrix elements at energy $e_{0}$. With this choice, the ansatz in Eq.(10b) is equivalent to
$\frac{\overline{\Psi_{ij}\Psi_{ji}}}{\overline{{|\Psi_{e^{+}e^{+}}|}}^{2}}\simeq e^{-\frac{\phi^{\prime\prime}(e^{+})}{4L}\omega_{ij}^{2}},\quad E^{+}=Le^{+}={E_{i}+E_{j}\over 2}\ .$ (25)
@@
Figure 2: ETH correlations between the initial state and the observable $A=\hat{\sigma}^{x}_{1}$. (a) Scaling with the system size of the ETH predictions in Eqs.(10). The red and black dashed lines indicate $(\dim\mathcal{H})^{-1}$ and $(\dim\mathcal{H})^{-2}$ respectively (b) Smoothed averages $\overline{A_{ij}\Psi_{ij}}$ describing the Fourier transform of the relaxation dynamics increasing system size. Smoothed quantities at energy density $e_{0}=0.10$ with $\tau=4$.
@@ 4 Discussion and Conclusions
Let us remark that our results describe a wide class of initial states, including for example products $|a\rangle_{A}|b\rangle_{B}$ of energy eigenstates $|a\rangle_{A}$ and $|b\rangle_{B}$ of subsystems $A$ and $B$, that have recently motivated studies of the eigenstate correlations [29, 30, 31, 32, 33, 34]. For local $1d$ Hamiltonians, these states have energy fluctuations $\Delta_{E}^{2}=O(1)$ in $N$. While obeying Eq.(4), they do not have the form of a large deviation (9) and their survival probability (15) is known to decay exponentially in time [48]. Nevertheless, this class of states is naturally included in our ansatz on the state-observable correlations in Eqs.(10), with a difference in Eq.(10b) which generically reads
$\overline{|\Psi_{ij}|^{2}}=e^{-\Phi(E^{+}+\omega/2)+\Phi(E^{+}-\omega/2)}/Z^{2}.$ (28)
These state are discussed in [45], where we verify numerically the general scaling with the system size of Eqs.(10), and comment on the relation with the literature.
Our work opens a series of perspectives. At long times, hydrodynamic modes are expected to play a dominant role in equilibration dynamics [54, 55, 56], and it would be valuable to investigate how hydrodynamic description can be incorporated into the non-equilibrium ETH. Additionally, one could explore how the current ETH framework applies to integrable systems that equilibrate to a generalized Gibbs ensemble [11, 42, 57, 58] or in the presence of many-body quantum scars [59, 60, 61, 62]. Further questions concern the natural generalization of our ansatz to cross-correlations between different states $c_{i}^{a}=\langle E_{i}|\psi_{a}\rangle$ for multiple choices of $a$ and $i$ and in the higher-order correlations between states and observables [16, 33, 34].
@@ A.2 A constraint on the ansatz
Let us now see how these constraints imply that $\Psi$ is a projector by proving Eqs. (14).
We start by evaluating the off-diagonal with $i\neq j$:
@@ Appendix B A bound on high frequency tail of $g(\omega)$
As a starting point we introduce
$Z(\beta)=\langle\psi|e^{-\beta H}|\psi\rangle,$ (37)
generalizing $Z(0)$ defined in (6).
@@
$C_{\Psi}^{\beta}(t)\equiv{\langle\psi|e^{-\beta H}A(t)|\psi\rangle\over Z(\beta)}.$ (39)

### v5

@@ 1 Introduction
One of the central questions is how to extend the ETH framework to describe non-equilibrium dynamics [2, 35, 36, 37, 38, 39, 40, 41]. In this work, we propose a statistical ansatz for the matrix elements of the projector on the initial out-of-equilibrium state $\Psi=\ket{\psi}\bra{\psi}$ | written in the eigenbasis of the Hamiltonian. Notably, the non-equilibrium dynamics are encoded in the correlations between the initial state and the observable’s off-diagonal matrix elements, which we describe in our framework. After introducing the ansatz and verifying its consistency, we discuss its implications for the relaxation dynamics towards equilibrium and numerically verify it in a non-integrable one-dimensional spin chain. The main novelty of this work stems from the interplay between the standard ETH ansatz for the operator and the one that we propose for the projector via the correlations that characterize their matrix elements and that allows us to describe the non-equilibrium dynamics beyond the steady value.
@@ 2.2 Ansatz
Diagonal pseudo-random $\tilde{R}_{ii}$ also have zero average, but the particular value of their variance may depend on the symmetry class of $\hat{H}$ (e.g. GOE or GUE), as it is the case for $R_{ii}$ in standard ETH [44, 19]. Given that $\Psi_{ii}$ is positive, $(1+\tilde{R}_{ii})\geq 0$. Since $\Psi_{ij}$ is a product of two quasi-random numbers, this implies various constraints on the joint properties of $\tilde{R}_{ij}$. Crucially, these variables are exponentially weakly correlated with $R_{ij}$ of the original ETH ansatz (1),
$\overline{R_{ij}\tilde{R}_{ji}}=g_{A,\Psi}(e^{+},\omega_{ij})\,e^{-S(e^{+})/2}\ ,\quad e^{+}=(E_{i}+E_{j})/(2N),$ (8)
where $g_{A,\Psi}(e^{+},\omega_{ij})$ is an order one smooth function of its variables, which describes the correlations crucial for non-equilibrium dynamics. Note that with our notations $e^{+}$ is the energy density between two eigenstates, while $e_{0}$, the energy density of the initial state in Eq.(4). The existence of such correlations between the off-diagonal products of $\Psi_{ij}A_{ji}$, can be shown by averaging over the (random) phases of the eigenvectors of $\hat{H}$, see the App.A.1 Note that similar cross-correlations also exist between the matrix elements of different observables, say $\hat{A}$ and $\hat{B}$ [45, 16]. Within our notations this leads to a smooth function $F_{AB}^{(n)}(\omega)=f_{A}(\omega)f_{B}(-\omega)g_{AB}(\omega)$ with $g_{AB}(\omega_{ij})=\overline{R_{ij}^{A}R^{B}_{ji}}$. .
@@
$\overline{|\Psi_{ij}|^{2}}$ $=\frac{e^{-N(\phi(e_{i})+\phi(e_{j}))}}{Z}^{2}$ (10b)
@@
where we made explicit the dependence on the system size $N$. In the case of states conforming to (4), Eq.(10b) simplifies to
$\overline{|\Psi_{ij}|^{2}}=\frac{e^{-2N\phi(e^{+})}}{Z^{2}}\,e^{-\frac{\phi^{\prime\prime}(e^{+})}{4N}\omega_{ij}^{2}}\ ,$ (11)
where one expands the energies $E_{i,j}=E^{+}\pm\omega_{ij}/2$ in Eq.(10b) around $E^{+}$ and uses the assumption of large deviation.
A similar term, $e^{-\frac{\phi^{\prime\prime}(e^{+})}{8N}\omega_{ij}^{2}}$, should appear also in Eq.(10c), however, in the limit of large $N$ this can be neglected because $f_{A}(e,\omega)$ is expected to decay at large frequencies at least as $\exp(-\beta/4\omega_{ij})$ [46] and only close-by eigenvectors contribute to quantum averages.
Finally, we note that the product $\Psi_{ij}A_{ij}$ has large fluctuations compared to the average (10c), see the Appendix.
@@ 2.3 Consistency checks
At the leading order in $N$, this implies (see the Appendix),
@@
2.4 Statistical distribution of the matrix elements
In the spirit of Berry conjecture [47], or by analogy to Random Matrix Theory, at leading order, one can expect individual coefficients $c_{i}$, rescaled by their typical value, to be Gaussian random numbers. For the distribution of the diagonal elements $\Psi_{ii}=|c_{i}|^{2}$, this leads to the well-known Porter-Thomas distribution for $x=\Psi_{ii}/\overline{\Psi_{ii}}$ [48, 49].
We apply the same argument to the distribution of the re-scaled off-diagonal matrix elements $z=\Psi_{ij}/\sqrt{\overline{\Psi_{ij}\Psi_{ji}}}=\tilde{R}_{ij}$, treating them as a product of two independent Gaussian random variables. In the thermodynamic limit, it yields the modified Bessel function
$P(z)=\frac{1}{\pi}{K_{0}(|z|)}\simeq\frac{1}{\sqrt{2\pi}}\frac{e^{-|z|}}{\sqrt{|z|}}\ ,$ (16)
where on the right-end side, we substituted the asymptotic expansion for large $z$.
2.4.1 Ansatz for the higher-order cross correlations
Note that Eq.(16) describes the distribution of the rescaled matrix elements $\tilde{R}_{ij}$, but it does not contradict the existence of their cross-correlations. This is analogous to the ETH matrix elements $R_{ij}=A_{ij}/\sqrt{\overline{A_{ij}A_{ji}}}$, which are known to follow Gaussian distribution [50, 51, 52], while, within RMT, the random variables $A_{ij}$ follow a large deviation function [53, 19] and different matrix elements $A_{ij}$ are correlated [15, 39, 21, 19]. In the case of $\Psi_{ij}$, in addition to (14), there is an infinite number of relations coming from the fact that it is a projector, see Appendix A.2.
Our main focus is on the correlations between $\Psi$ and the observables, such as the ones in Eq. (10c). For a multiple product with the distinct indices $i_{1}\neq i_{2}\neq i_{n}$, one can write
$\overline{\Psi_{i_{1}i_{2}}A_{i_{2}i_{3}}...A_{i_{n}i_{1}}}=e^{-(n-1)S(e_{+})}\frac{e^{-\Phi(E_{+})}}{Z}h_{e^{+}}^{(n)}(\vec{\omega})\ ,$ (17)
while contributions with the repeated indices factorize, i.e.
$\overline{\Psi_{i_{1}i_{2}}...A_{i_{k-1}i_{1}}A_{i_{1}i_{k+1}}\dots A_{i_{n}i_{1}}}=\overline{\Psi_{i_{1}i_{2}}...A_{i_{k-1}i_{1}}}\cdot\overline{A_{i_{1}i_{k+1}}\dots A_{i_{n}i_{1}}}+O({\mathcal{D}}^{-1})\ .$ (18)
Here $h_{e^{+}}^{(n)}(\vec{\omega})$ is a smooth function of its arguments: $e^{+}=(E_{i_{1}}+\dots E_{i_{n}})/Nn$ and $\vec{\omega}=(E_{i_{1}}-E_{i_{2}}$ $,\dots,E_{i_{n}}-E_{i_{n-1}})$. This ansatz encodes the out-of-equilibrium multi-time dynamics such as
$\bra{\psi}\hat{A}(t_{1})\hat{A}(t_{2})\dots\hat{A}(t_{n-1})\ket{\psi}$ and also ensures that the ansatz on the initial state is stable against perturbations.
@@ 2.5 Implications for the dynamics
By substituting sums with integrals, neglecting spectral correlations, and using the out-of-equilibrium ETH ansatz in Eq.(11), at the leading order in $N$, one finds
@@
where we used the definition of the energy variance in Eq.(12), $\Delta_{E_{0}}=1/\sqrt{\Phi^{\prime\prime}(E_{0})-S^{\prime\prime}(E_{0})}$ $=\delta_{e_{0}}\sqrt{N}$. Thus the large deviation ansatz in Eq.(6) is consistent with the Gaussian decay, controlled by the energy variance of the initial state, in agreement with the literature on global quenches, see e.g. Ref. [55]. Dynamical behavior, different from Eq.(20), e.g. including an exponential decay, is known to arise from the initial states which are different from Eq.(4) [54, 58, 59, 55, 56, 60, 61], as discussed below.
@@
Similarly to $f_{A}^{2}(\omega)$, which has to decay exponentially at large $\omega$ in $D\geq 2$ (superexponentially in 1D), high-frequency tail of $g_{\Psi,A}(\omega)$ has to be exponentially suppressed for states $\Psi$ associated with local perturbations (see the Appendix).
@@ 3 Numerical results
Figure 1: Out-of-Equilibrium ETH of the fully polarized initial state ${|\psi\rangle=|\downarrow_{z}\dots\downarrow_{z}\rangle}$. (a) The rescaled diagonal ensemble as a function of the energy density for different system sizes $L=12,14,16,18$. In the inset, the initial energy $E_{0}=\langle\psi|\hat{H}|\psi\rangle$ and the energy fluctuations $(\Delta E_{0})^{2}=\langle\psi|\hat{H}^{2}|\psi\rangle-E_{0}^{2}$ are plotted as a function of the inverse of the system size $1/L$. (b) The off-diagonal matrix elements of the projector $\Psi_{ij}$ as a function of the energy difference $\omega_{ij}=E_{i}-E_{j}$ rescaled by $1/\sqrt{L}$. In both panels, the blue dots correspond to individual overlaps for $L=16$. The smoothing parameter is $\tau=4$. In the inset of panel (b), the numerical fidelity decay up to $L=22$ is compared with the prediction of Eq.(20) (dashed line) without any fitting parameters.
@@
First, we establish that the initial states (27) are consistent with the ansatz in Eq.(6). In Fig. 1a we plot the diagonal ensemble for different length sizes $L=12,14,16,18$, showing that it obeys the large deviation prediction $\overline{\Psi_{ii}}=\frac{e^{-L\phi(E_{i}/L)}}{Z}$. This is confirmed by the inset, where we plot the scaling of the initial energy $E_{0}=e_{0}L$ and variance $\Delta_{E_{0}}/E_{0}=\delta_{e_{0}}/e_{0}\sqrt{L}$, c.f. Eq.(12). From a fit of the data, we extract the dimensionless values $e_{0}=0.10$, $\delta_{e_{0}}=1.12$. In panel (b), we study the fluctuations of the out-of-equilibrium ETH functions in the frequency domain [cf. Eq.(10b)]. To address the dependence on $L$, we re-scale $\overline{|\Psi_{ij}|^{2}}$ by the diagonal matrix elements at energy $e_{0}$ defined in Eq.(4). For the state under consideration, the ansatz in Eq.(10b) is given by Eq.(11) and it is given by
$\frac{\overline{\Psi_{ij}\Psi_{ji}}}{\overline{{|\Psi_{e_{0}e_{0}}|}}^{2}}\simeq e^{-\frac{\phi^{\prime\prime}(e_{0})}{4L}\omega_{ij}^{2}}\ .$ (29)
@@
Figure 2: Numerical distributions of the projector’s matrix elements $P(\tilde{R}_{ij})$ for $L=18$ superimposed with the theoretical predictions. (a) Diagonal matrix elements $x=\Psi_{ii}/\overline{\Psi_{ii}}$ (blue) and the Porter-Thomas distribution in Eq.(30) (orange). (b) Off-diagonal matrix elements $z=\Psi_{ij}/\overline{\Psi_{e_{0}e_{0}}}$ (blue) superimposed with the modified Bessel function Eq. (16) (orange). The data is for the mean energy density $e_{0}=0.10$ smoothed with $\tau=4$.
We show the distribution of matrix elements $\tilde{R}_{ij}$ in Fig. 2. In panel (a), we show rescaled diagonal matrix elements $x=\Psi_{ii}/\overline{\Psi_{ii}}$, $\Psi_{ii}=|c_{i}|^{2}$. The data is in agreement with the Porter-Thomas distribution [48], which is expected for systems with time-reversal symmetry, falling in the orthogonal ensemble (OE) universality class,
$P_{\rm OE}(x)=\frac{e^{-x/2}}{\sqrt{2\pi x}}\ .$ (30)
In panel (b), we report the data for the off-diagonal matrix elements, rescaled as $z=\Psi_{ij}/\overline{\Psi_{e_{0}e_{0}}}$, where $\Psi_{e_{0}e_{0}}$ is defined in Eq. (29). Our numerical results are in good agreement with Eq. (16), without any fitting parameter.
Figure 3: ETH correlations between the initial state and the observable $A=\hat{\sigma}^{x}_{1}$. (a) Scaling with the system size of the ETH predictions in Eqs.(10). The red and black dashed lines indicate $(\dim\mathcal{H})^{-1}$ and $(\dim\mathcal{H})^{-2}$ respectively (b) Smoothed averages $\overline{A_{ij}\Psi_{ij}}$ describing the Fourier transform of the relaxation dynamics increasing system size. The data is for the mean energy density $e_{0}=0.10$ smoothed with $\tau=4$.
@@ 4 Discussion and Conclusions
Let us remark that our results describe a wide class of initial states, including for example products $|a\rangle_{A}|b\rangle_{B}$ of energy eigenstates $|a\rangle_{A}$ and $|b\rangle_{B}$ of subsystems $A$ and $B$, that have recently motivated studies of the eigenstate correlations [29, 30, 31, 32, 33, 34]. For local $1d$ Hamiltonians, these states have energy fluctuations $\Delta_{E}^{2}=O(1)$ in $N$. While obeying Eq.(4), they do not have the form of a large deviation (9) and their survival probability (19) is known to decay exponentially in time [55]. Nevertheless, this class of states is naturally included in our ansatz on the state-observable correlations in Eqs.(10), where Eq.(10b) generically reads
$\overline{|\Psi_{ij}|^{2}}=e^{-\Phi(E^{+}+\omega/2)-\Phi(E^{+}-\omega/2)}/Z^{2}.$ (33)
These states are discussed in the Appendix, where we verify numerically the general scaling with the system size of Eqs.(10), and comment on the relation with the literature.
Our work opens a series of perspectives. At long times, hydrodynamic modes are expected to play a dominant role in equilibration dynamics [62, 63, 64], and it would be valuable to investigate how hydrodynamic description can be incorporated into the non-equilibrium ETH. Additionally, one could explore how the current ETH framework applies to integrable systems that equilibrate to a generalized Gibbs ensemble [11, 42, 65, 66] or in the presence of many-body quantum scars [67, 68, 69, 70].
@@ A.2 A constraint on the ansatz
Let us now see how these constraints imply that $\Psi$ is a projector by proving Eqs. (15). We start by evaluating the off-diagonal with $i\neq j$:
@@
In addition to (36) and (37), there is an infinite number of higher-order constrains coming from the fact that $\Psi_{ij}$ is a projector. Say, we find at order $3$,
$\tilde{R}_{ij}\tilde{R}_{jk}\tilde{R}_{ki}=\tilde{R}_{ik}(1+\tilde{R}_{jj})\tilde{R}_{ki}=(1+\tilde{R}_{ii})(1+\tilde{R}_{jj})(1+\tilde{R}_{kk}),$ (41)
for $i\neq j\neq k$ as well as for $i\neq j\neq k\neq l$ at order $4$,
$\tilde{R}_{ij}\tilde{R}_{jk}\tilde{R}_{kl}\tilde{R}_{li}=\tilde{R}_{ik}(1+\tilde{R}_{jj})\tilde{R}_{ki}(1+\tilde{R}_{ll})=(1+\tilde{R}_{ii})(1+\tilde{R}_{jj})(1+\tilde{R}_{kk})(1+\tilde{R}_{ll}).$ (42)
The list of such constraints continues to include higher orders of $\tilde{R}$.
@@ Appendix B A bound on high frequency tail of $g(\omega)$
As a starting point, we introduce
$Z(\beta)=\sum_{i}e^{-\Phi(E_{i})-\beta E_{i}}$ (44)
such that $Z=Z(0)$ matches the definition of $Z$ in (6). It is also convenient to introduce the ratio
${\cal Z}(\beta)=\langle\psi|e^{-\beta H}|\psi\rangle={Z(\beta)\over Z}.$ (45)
@@
$C_{\Psi}^{\beta}(t)\equiv{\langle\psi|e^{-\beta H}A(t)|\psi\rangle\over{{\cal Z}}(\beta)}.$ (47)